# Unsupervised Model Learning for Quantum Impurity Systems

### v1

@@ II Effective Hamiltonians
$+\hat{H}^{\text{bath}}_{\rm disc}+V\sum_{\sigma}(\hat{d}_{\sigma}^{\dagger}\hat{c}_{0\sigma}+\hat{c}_{0\sigma}^{\dagger}\hat{d}_{\sigma})\;.$ (23)
It is reasonable to assume that the potential energy is lower than the kinetic energy in realistic QIMs, meaning we choose sAIM parameters with $U_{d}/D<1$ to keep the bandwidth the largest energy scale in the system. However, the sAIM transforms under RG such that $U_{d}/D\rightarrow U^{\prime}_{d}/D^{\prime}\gg 1$ and thus the interaction becomes the dominant energy scale in the system, this is known as the local moment (LM) FP. Under these circumstances one can perform a second order perturbation theory in the impurity-bath hybridization $V$ to obtain the famous Kondo model
$\hat{H}^{\text{K}}_{\rm eff}=\hat{H}^{\text{bath}}+J\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\;,$ (24)
where $\hat{\mathbf{S}}_{0}=\tfrac{1}{2}\sum_{\sigma,\sigma^{\prime}}\boldsymbol{\sigma}_{\sigma\sigma^{\prime}}c_{0\sigma}^{\dagger}c_{0\sigma^{\prime}}$ is the conduction electron spin density at the impurity position. The Kondo model contains only terms that are RG-relevant to the flow of the sAIM and it encapsulates all its physics even for further scaling, away from the LM FP, this is the SWT. The SWT can also be applied to the non-renormalized theory, for which it yields $J_{\rm SW}=8V^{2}/U_{d}$. However, this does not account for bandwidth renormalization $D\to D_{\rm eff}$. An approximate form of the bandwidth renormalization in the small $U_{d}$ limit was found by Haldane [63], and discussed further by Krishnamurthy et al in [50], but still does not include subleading corrections from finite $J$. Recently, the authors determined the precise form numerically for all $U_{d}$ and $J_{\rm SW}$ using an alternative machine learning scheme based on optimization with respect to the partition function [47]. There it was shown that while the $J$ of the effective Kondo model, Eq. 24, can be tuned to reproduce the correct low-energy physics and Kondo scale $T_{\rm K}$, a coinciding free energy does not imply that all local observables are simultaneously matched between bare and effective model since the RG-flow of the bare and effective models is not the same. Therefore, the KLD is a somewhat misleading measure of distinguishability as its minimum requires local observables to match, but matching observables does not imply a matching free energy or matching low-energy scales. The KLD is only then a good measure of distinguishability if the bare and effective model have the same RG-flow, which is not the case when only RG-relevant terms are included in the effective model. Therefore, in this chapter we go beyond the minimal models, like the Kondo model and introduce minimally constrained effective models. These models are completely general, determined only by the symmetries of the bare model and thus, also allow to include RG-marginal or -irrelevant terms. We show in the following sections that such an effective model has a RG-flow that is comparable to the bare model and it is hence capable of reproducing a special set of local observables as well as the universal low-energy Kondo physics.
The RG-analysis that yields the minimal model also provides insights in the reduction of degrees of freedom of the impurity; for sufficiently low temperatures ($T\lesssim U_{d}$) the impurity turns from a spin-full, fermionic orbital into a spin-1/2 magnetic moment. The only aspect that we adopt from the minimal model in the minimally constrained model is the effective impurity Fock space $\mathcal{H}^{\rm imp}_{\rm eff}$ that replaces the bare Fock space $\mathcal{H}^{\rm imp}_{\rm bare}$. To meaningfully define an impurity Hamiltonian, a Fock space $\mathcal{H}^{\rm imp}_{\rm eff}$ of a spin is not enough, at least one bath site has to be added to capture the impurity-bath coupling:
$\mathcal{H}^{\rm bath}_{\rm 0}\otimes\mathcal{H}^{\rm imp}_{\rm eff}=\mathcal{H}^{\rm 1e-imp}_{\rm eff}\;,$
we refer to this as single extended impurity ($\rm 1e-imp$). To systematically generalise this we consider the semi-infinite Wilson chain Eq. 13. This allows the construction of an $n$-extended impurity
$\mathcal{H}^{n\rm e-imp}_{\rm eff}=\mathcal{H}^{\rm bath}_{\rm n}\otimes\mathcal{H}^{\rm bath}_{\rm n-1}\otimes\dots\otimes\mathcal{H}^{\rm bath}_{\rm 0}\otimes\mathcal{H}^{\rm imp}_{\rm eff}\;,$
where $\mathcal{H}^{\rm bath}_{i}$ corresponds to the Fock space of Wilson site $i$. Note that every site that is absorbed in the extended impurity is removed from the bath
$\hat{H}^{\rm bath}_{\rm disc}\mapsto\hat{\tilde{H}}^{\rm bath}_{\rm disc}=\sum_{\sigma}\sum^{\infty}_{i=n}\dots\;.$
The enlarged Fock space permits the inclusion of high order exchange terms with the bath and, due to the nature of the Wilson chain, RG-marginal and -irrelevant terms. It depends on the precise FP one is considering which eigenvalues the $i^{\rm th}$ bath site operator $\hat{c}^{(\dagger)}_{\sigma i}$ has in terms of the linearized RG transformation $\mathcal{R}^{*}_{l}$ (Eq. 21), but impurity operators have in general eigenvalues with $|\lambda^{*}|>1$ and bath operators $|\lambda^{*}|<1$, with $|\lambda^{*}|$ decreasing for increasing $i$. Thus, including bath sites allows one to construct effective interactions $\hat{h}_{i}$ that go beyond the FP point physics and are closer to the bare model in terms of the RG-flow. Thereby we systematically improve the quality of our effective model.
Following Ref. [50] we construct the minimally constrained effective Hamiltonian on the $n$-extended impurity Fock space, with the only constraints being the symmetries of the bare model $\{\rm S_{1},S_{2},\dots\}$. In the following we only consider models invariant under $U(1)\equiv\textrm{Q}$ and $U(2)\equiv\textrm{S}$ S as they are most commonly encountered in quantum impurity models and in fact also in the sAIM, however our approach can be generalised to more symmetries. Since representations of Q and S on $\mathcal{H}^{\rm 1e-imp}_{\rm eff}$ commute we can construct a basis on the effective Fock space and label it with the quantum numbers $S$ for the non-abelian spin S quantum number, $s_{z}$ for the abelian magnetic quantum number of the Cartan subalgebra of S, and $Q$ for the abelian charge quantum number
where $m$ is the multiplet index, as the same representation of $S$ and $Q$ can be repeated several times in the same Fock space. The multiplets are not unique: every linear combination of a multiplet of some collection of quantum numbers is again a multiplet with the same quantum numbers, hence multiplets introduce a gauge degree of freedom in the choice of the basis. Therefore, it is not possible to compare degenerate multiplets between bare and effective models. We show with the example of the sAIM that we can eliminate this ambiguity and enable meaningful analogies between bare and effective states. For now, assume we have chosen a multiplet gauge and thus, we can construct projectors on the quantum number subspaces
$\hat{P}_{QS}=\sum_{m}\ket{Q,S;m}\bra{Q,S;m}.\;,$ (26)
where omitting $s_{z}$ implies that we have summed over all $s_{z}=-S,-S+1\dots,S$. These projectors form the operator basis that we use to construct the effective impurity Hamiltonian
$\hat{H}^{n\rm-imp}_{\rm eff}(\boldsymbol{\theta})=\sum_{Q,S}\theta_{QS}\hat{P}_{QS}\;,$ (27)
and according to our previous discussion, this leads to a convex KLD as the projectors clearly all mutually commute. This is the ideal case in terms of optimisation, but it is not the most general $SU(2)\times U(1)$ Hamiltonian on $\mathcal{H}^{n\rm e-imp}_{\rm eff}$, since different multiplets with the same QNs are treated independently in $\hat{H}_{\rm eff}$, but may also be connected. With the example of the sAIM we make $\hat{H}_{\rm eff}$ more general in the next section at the cost of the convexity of the KLD. However, here we first stick with Eq. 27. One of the advantages of the projectors $\hat{P}_{QS}$ is that ${\hat{\Omega}_{\rm ad}^{\dagger}\cdot\hat{P}_{QS}\cdot\hat{\Omega}_{\rm ad}}$ is trivial to compute and it makes the observables directly comparable between bare and effective model. This is because bare and effective model have the same symmetries by construction.
An important property of the projectors is that the sum over all projectors gives the identity on the effective Fock space $\mathcal{H}^{n\rm e-imp}_{\rm eff}$
$\hat{\mathbbm{1}}_{\rm eff}=\sum_{Q,S}\hat{P}_{QS}\;,$ (28)
from which follows that $\langle\hat{\mathbbm{1}}_{\rm eff}\rangle_{\rm eff}=1$. But the sum over all projectors is clearly not the identity on the bare Fock space and therefore $\langle\hat{\Omega}^{\dagger}_{\rm ad}\cdot\hat{\mathbbm{1}}_{\rm eff}\cdot\hat{\Omega}_{\rm ad}\rangle_{\rm bare}\leq 1$. From this it follows that it is impossible to demand
for all $Q$ and $S$ combinations at the same time, as this would amount to demanding
$\langle\hat{\mathbbm{1}}_{\rm eff}\rangle_{\rm eff}\overset{!}{=}\langle\hat{\Omega}^{\dagger}_{\rm ad}\cdot\hat{\mathbbm{1}}_{\rm eff}\cdot\hat{\Omega}_{\rm ad}\rangle_{\rm bare}\;,$ (30)
which cannot be satisfied for a meaningful reduction in complexity of $\hat{H}_{\rm eff}$ to $\hat{H}_{\rm bare}$. Therefore at least one projector (or a multiplet within a projector) must be omitted in the Hamiltonian $\theta_{QS}=0$, this choice is not arbitrary and an ideal choice can be made, as we show in Sec. IV.
III Single and double extended effective impurity models
The simplest effective Hamiltonians for the AIM are defined on a single extended (1e-imp) Fock space and the prime example is the Kondo model. However the Kondo model can only either reproduce the local observables or the universal temperature scale $T_{K}$ and not both, as shown recently by the authors [47]. The minimally constrained models get increasingly capable of reproducing both by increasing the number of bath sites included in the extended impurity.
The construction of the 1e-imp minimally constrained model begins with creating the basis of the $\mathcal{H}^{1\rm e-imp}_{\rm eff}$ Fock space, which is straight forward as $M=1$ for all QN combinations. The basis states are then turned into projectors associated with Q and S

### v2

@@ II.3 Schrieffer-Wolff revisited
$\begin{split}\hat{H}^{\rm AIM{}}_{\text{bare}}=&\hat{H}^{\text{bath}}_{\rm disc}+\tfrac{1}{2}U_{d}(\hat{n}^{d}-1)^{2}\\ &+V\sum_{\sigma}(\hat{d}_{\sigma}^{\dagger}\hat{c}_{0\sigma}^{\phantom{\dagger}}+\hat{c}_{0\sigma}^{\dagger}\hat{d}_{\sigma}^{\phantom{\dagger}})\;,\end{split}$ (24)
where $\hat{n}^{d}=\sum_{\sigma}\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$. As defined, the model possesses particle-hole symmetry and is hence at half filling (the ground state has impurity charge $Q=1$). For $U_{d}\gg\Gamma$, with $\Gamma=\pi V^{2}/2D$ the impurity-bath hybridization, local moment formation on the impurity is expected. Under these circumstances, second order perturbation theory in the impurity-bath coupling $V$ yields the Kondo model,
$\hat{H}^{\text{K}}_{\rm eff}=\hat{H}^{\text{bath}}_{\rm disc}+J\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\;,$ (25)
where $\hat{\mathbf{S}}_{0}=\tfrac{1}{2}\sum_{\sigma,\sigma^{\prime}}\boldsymbol{\sigma}_{\sigma\sigma^{\prime}}c_{0\sigma}^{\dagger}c_{0\sigma^{\prime}}^{\phantom{\dagger}}$ is the conduction electron spin density at the impurity position. This is the famous Schrieffer-Wolff transformation (SWT) [66, 36]. The Kondo model is as such the minimal effective model for the AIM, in the sense defined in the previous section.
The second-order perturbative calculation yields an estimate for the spin-exchange coupling, $J_{\rm SW}=8V^{2}/U_{d}$, which is naturally a rather crude approximation at finite $U_{d}/\Gamma$. Indeed, as discussed in Ref. [48], physical properties such as the thermodynamic entropy and Kondo temperature of the effective model Eq. 25 computed with $J_{\rm SW}$ do not match well those of the bare AIM Eq. 24 unless $U_{d}\gg D$. In fact, even when the SWT is carried out to infinite order in the impurity-bath hybridization [67], the predicted value of $J$ still does not give accurate results for low-energy properties [48]. The reason for this is well known: the SWT does not account for the bandwidth renormalization $D\to D_{\rm eff}$. In the small $U_{d}/D$ limit, Haldane found a perturbative estimate for this effect [68], with the intuitive result that $D_{\rm eff}\sim U_{d}$. This was put on a more concrete footing by Krishnamurthy et al in Ref. [51] using RG arguments. Only recently in Ref. [48] was the precise form determined numerically by non-perturbative methods for all $U_{d}$ and $V$. The conclusion is that even with this highly simplified minimal effective model, the free energy and hence low-temperature thermodynamics can be faithfully reproduced – provided the correct effective coupling $J$ is found.
II.4 $\mathcal{F}$-learning for the Kondo model
Since the AIM and the Kondo model are connected by the RG flow of the AIM, the low-temperature thermodynamics of the AIM, such as the impurity entropy flow, magnetic susceptibility, and the Kondo temperature $T_{K}$, can be exactly reproduced in an effective Kondo model with a properly chosen coupling constant $J$, provided $T_{K}\ll U_{d}$. As shown in Ref. [48], this ‘true’ $J$ can be found numerically exactly by solving the AIM and Kondo models non-perturbatively using e.g. NRG, and matching their free energies
$\mathcal{F}^{\rm eff}(J)=\mathcal{F}^{\rm bare}(U_{d},V)\;,$ (26)
which according to Eq. 20 also matches their RG flows. This should be done at sufficiently low temperatures $T\ll U_{d}$ where the RG flows of the two models are expected to coincide. We refer to this model learning technique as $\mathcal{F}$-learning. This is a powerful method if one is interested in reproducing the low-temperature thermodynamics or low-frequency dynamical correlation functions of some bare model using a simpler effective model – provided that both bare and effective models can be solved accurately down to a temperature $T\ll U_{d}$. However, for complex impurity problems, such a direct solution of the bare model might not be feasible. Furthermore, a coinciding free energy in bare and effective models does not guarantee that all local thermodynamic observables will agree, as discussed in the next subsection. If such observables are the desired outcome from an effective model, $\mathcal{F}$-learning alone is not the best choice.
II.5 Renormalization of observables
Thermodynamic observables may themselves flow under RG, and the full path taken, from high to low energies, affects the result. Since the effective model is defined on an energy scale below some cutoff, it cannot capture renormalization effects in the bare model at energies above this cutoff. The important implication of this is that we cannot expect even a optimal minimal effective model to capture simultaneously low-energy scales such as the Kondo temperature $T_{K}$, as well as the behavior of thermal expectation values. The former is controlled by the low-temperature free energy, whereas the latter depends on the entire RG flow.
We demonstrate this point explicitly for the simplest example of the mapping from the AIM to the Kondo model. Using $\mathcal{F}$-learning [48] we extracted the effective $J$ of the Kondo model that reproduces exactly the low-temperature thermodynamics and Kondo temperature $T_{K}$, for reference AIM systems with different bare $U_{d}$ (keeping $V^{2}/U_{d}$ constant and setting $D\equiv 1$). For each bare AIM and corresponding $\mathcal{F}$-optimized Kondo model, we computed with NRG the zero-temperature expectation value of the spin-spin correlator $\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle$. The results are shown in the main panel of Fig. 2, demonstrating markedly different renormalization of this observable in bare and effective models, especially at small $U_{d}$. The ratio of the observables calculated in bare and effective models, shown in the inset, characterizes the different degree of renormalization in the two models along the RG flow. Only for $U_{d}\gtrsim 100D$ do the observables agree; this is consistent with the fact that the perturbative SWT becomes asymptotically exact in the large $U_{d}/\Gamma$ limit.
Figure 2: Impurity-bath spin-spin expectation value ${\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle}$ computed with NRG at $T=0$ for the AIM Eq. 24 (blue line) compared with the corresponding value in the Kondo model Eq. 25, optimized by $\mathcal{F}$-learning (green). The inset shows their ratio, which captures the different renormalizations along the two paths to the same low-energy fixed point. Plotted as a function of $U_{d}$ for $0.38V^{2}/U_{d}=$ 0.3 and $D=1$.
One can understand this result using the framework of Wilsonian RG described above. It is instructive to interpret the RG procedure as a reparameterization of the free energy, $\mathcal{F}$. In this picture, the renormalization is achieved by replacing the original coordinate system $\mathbf{C}$ with a new coordinate system $\mathbf{C}^{\prime}$ that leaves $\mathcal{F}$ invariant, as required by Eq. 20. Thus we have $\mathbf{C}\rightarrow\mathbf{C}^{\prime}(\mathbf{C})$ and thermodynamic observables can be interpreted as covariant tensors that can be transformed as such between coordinate systems [63]. For the effective model Eq. 3, we may therefore write,
$\langle\hat{h}_{i}\rangle_{\rm eff}$ $=\frac{\partial}{\partial\theta_{i}}\mathcal{F}_{\rm eff}(\boldsymbol{\theta})$ (27)
$=\sum_{j}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\frac{\partial}{\partial\theta^{*}_{j}}\mathcal{F}_{\rm eff}(\boldsymbol{\theta}^{*})=\sum_{j}\chi^{*}_{\theta_{j}}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\;,$
with $\boldsymbol{\theta}^{*}$ being the renormalized effective model parameters at the low-energy FP, and where $\chi^{*}_{\theta_{j}}$ is the expectation value of the operator $\hat{h}_{j}$ evaluated in the FP Hamiltonian. The matrix of derivatives $\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}$ contains information about the RG flow to the FP, and can in principle be obtained from the $\beta$-function, Eq. 21. With this information we can make a connection between the renormalized value of the observable and its FP value. The desired observable is found to be a linear combination of all possible FP expectation values.
Taking again the example of the AIM to Kondo model mapping, the low-energy stable FP (of both models) is of course the strong coupling (SC) FP [31, 51]. In the effective Kondo model, the SC FP Hamiltonian is characterized by $J^{*}\to\infty$. Therefore ${\chi^{*}_{J}\equiv\partial_{J^{*}}\mathcal{F}_{\rm eff}(J^{*})={\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle}^{*}}$ evaluated at the SC FP takes the value,
$\chi^{*}_{J}=-3/4\;,$ (28)
which as such embodies the spin-singlet formed between the impurity and bath. However the actual value of the correlator as measured in the Kondo model, ${\chi_{J}\equiv\partial_{J}\mathcal{F}_{\rm eff}(J)}$, carries information about the RG flow,
$\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle=-\frac{3}{4}\frac{\partial J^{*}}{\partial J}\;.$ (29)
It is tempting to use perturbative scaling techniques [69] to evaluate the nontrivial factor $\tfrac{\partial J^{*}}{\partial J}$, but we find that these yield rather poor approximations to the true value of the correlator obtained by NRG.
To understand the flow of the same observables $\langle\hat{h}_{i}\rangle$ in the bare model, we have to add $\hat{h}_{i}$ to the bare model as a source term. This is because the effective interactions $\hat{h}_{i}$ do not typically appear in the original bare model. Thus we may write,
$\langle\hat{h}_{i}\rangle_{\rm bare}$ $=\frac{\partial}{\partial\theta_{i}}\mathcal{F}_{\rm bare}(\mathbf{C};\theta_{i})\Big{|}_{\theta_{i}=0}$ (30)
$=\left[\frac{\partial\theta^{*}_{i}}{\partial\theta_{i}}\frac{\partial}{\partial\theta^{*}_{i}}+\sum_{j}\frac{\partial C^{*}_{j}}{\partial\theta_{i}}\frac{\partial}{\partial C^{*}_{j}}\right]\mathcal{F}_{\rm bare}(\mathbf{C}^{*};\theta^{*}_{i})\Big{\rvert}_{\theta_{i}=0}$
where $\mathbf{C}^{*}$ are the bare model FP parameters, and $\chi^{*}_{\theta_{i}}$ or $\chi^{*}_{C_{j}}$ are expectation values evaluated in the FP Hamiltonian of the bare model.
In the low-energy limit, both bare and effective models share the same FP, by construction. Therefore in the bare model we may write the FP parameters $\theta_{j}^{*}\equiv\theta^{*}_{j}({\mathbf{C}})$ as functions of the bare model parameters. From this it follows that
$\langle\hat{h}_{i}\rangle_{\rm bare}=\sum_{j}\chi^{*}_{\theta_{j}}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\Big{|}_{\theta_{i}=0}\;,$ (31)
where $\chi_{\theta_{j}^{*}}$ are the FP expectation values of the effective model operators $\hat{h}_{j}$. Since the FP is the same, these expectation values are the same in bare and effective models. For the AIM to Kondo mapping, we find,
$\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle_{\rm eff}\leq\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle_{\rm bare}\;,$ (32)
as demonstrated numerically in Fig. 2. This implies that,
$\left[\frac{\partial J^{*}}{\partial J}\right]_{\rm bare}\Big{/}\left[\frac{\partial J^{*}}{\partial J}\Big{|}_{J=0}\right]_{\rm eff}\leq 1\;,$ (33)
where equality is only approached in the large $U_{d}$ limit of the AIM where the RG flow is equivalent to that of the Kondo model.
The above has consequences for automated model learning. When optimizing an effective model with UML using GD, Eq. 10, we should match the full set of observables $\langle\hat{h}_{i}\rangle_{\rm bare}=\langle\hat{h}_{i}\rangle_{\rm eff}$ for all effective model operators $\hat{h}_{i}$. This in turn implies that we enforce,
$\Bigg{[}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\Bigg{]}_{\rm eff}=\Bigg{[}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\Big{|}_{\theta_{i}=0}\Bigg{]}_{\rm bare}\;.$ (34)
However, the RG flow to reach the low-energy FP is not the same in bare and effective models since the latter typically has fewer FPs. Therefore enforcing Eq. 34 will give a solution for which the free energy of the effective model is in general not the same as that of the bare model. The low-temperature thermodynamic entropy, Kondo scales etc, are therefore only in good agreement if the matched observables do not flow strongly. For the Kondo model, we see that $\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle$ is in general renormalized rather differently compared with the AIM.
Our conclusion is that the KLD Eq. 9 is only a good measure of Hamiltonian distinguishability if the bare and effective models have similar RG flows. A minimal effective model may therefore be poorly suited to this kind of optimization.
The question addressed in the next section is: can more general effective model structures in principle capture the Kondo scale and thermal expectation values simultaneously, at least approximately? If so, can UML be used to automate the optimization while remaining computationally tractable?
III Minimally-constrained
    effective models
We now go beyond minimal models, like the Kondo model, and introduce a family of minimally constrained (MC) effective models. With a given choice for the reduced Fock space of the effective model, MC models are characterized by a completely general operator structure, constrained only to respect the symmetries of the bare model. This allows for the inclusion and parameterization of all possible RG relevant, marginal and irrelevant operators. We show in the following sections that such an e ffective model can much more accurately capture the physics of the bare model. Because its RG flow is more comparable with the bare model, MC models are better able to capture both the behavior of local observables as well as the Kondo temperature and associated universal low-energy Kondo physics. This also means that the UML strategy can be used to find good effective models, when the starting point is an MC model.
To construct an MC model, we first must decide on the reduced Fock space in which the effective model lives. In this regard, we note that by replacing the impurity part of the bare Hamiltonian by operators spanning only the ground state manifold, we also implicitly change the nature of the impurity-bath hybridization term. Consider for example the classic mapping of the AIM to the Kondo model. The impurity degrees of freedom $\hat{d}_{\sigma}$ are mapped to a spin-$\tfrac{1}{2}$ operator $\hat{\mathbf{S}}_{d}$, but the impurity-bath tunneling $V\sum_{\sigma}(\hat{d}_{\sigma}^{\dagger}\hat{c}_{0\sigma}^{\phantom{\dagger}}+{\rm H.c.})$ in Eq. 24 is also replaced by a spin-spin exchange coupling $J\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$ in Eq. 25.
An alternative way to formulate the mapping is to take an extended view on what constitutes the ‘impurity’ to include also the local bath orbitals to which the bare impurity degrees of freedom couple. In so doing, we also redefine our bath to exclude these local orbitals. We thus repartition the bare model as $\hat{H}_{\rm bare}=\hat{\tilde{H}}^{\rm imp}_{\rm bare}+\hat{H}^{\rm bath}_{1}+\hat{H}^{\rm hyb}_{1}$ where our extended impurity is now given by,
$\hat{\tilde{H}}^{\rm imp}_{\rm bare}=\hat{H}_{\rm bare}^{\rm imp}+\hat{H}_{\rm bare}^{\rm hyb}\;.$ (35)
The extended impurity Fock space is thus,
~$\tilde{\mathcal{H}}^{\rm imp}_{\rm bare}=\mathcal{H}^{\rm imp}_{\rm bare}\otimes\mathcal{H}^{\rm bath}_{\rm 0}\;.$ (36)
The bath and hybridization terms are correspondingly redefined,
$\hat{H}^{\rm bath}_{N}$ $=\sum^{\infty}_{n=N}\sum_{\sigma}t_{n}^{\phantom{\dagger}}(\hat{c}^{\dagger}_{n\sigma}\hat{c}_{n+1\sigma}^{\phantom{\dagger}}+\hat{c}^{\dagger}_{n+1\sigma}\hat{c}_{n\sigma}^{\phantom{\dagger}})\;,$ (37)
$\hat{H}^{\rm hyb}_{N}$ $=\sum_{\sigma}t_{N-1}^{\phantom{\dagger}}(\hat{c}^{\dagger}_{N-1\sigma}\hat{c}_{N\sigma}^{\phantom{\dagger}}+\hat{c}^{\dagger}_{N\sigma}\hat{c}_{N-1\sigma}^{\phantom{\dagger}})\;,$ (38)
where $N=1$ indicates the new starting site index of the redefined bath chain. This has the advantage that the new bath and hybridization terms are always the same in bare and effective models. We have emphasized this by now removing the ‘bare’ label from these terms.
The MC effective model is similarly structured,
$\hat{H}^{\rm MC}_{\rm eff}(\boldsymbol{\theta})=\hat{\tilde{H}}^{\rm imp}_{\rm eff}(\boldsymbol{\theta})+\hat{H}^{\rm bath}_{1}+\hat{H}^{\rm hyb}_{1}\;,$ (39)
where, $\tilde{\tilde{H}}^{\rm imp}_{\rm eff}(\boldsymbol{\theta})=\sum_{i}\theta_{i}\hat{h}_{i}$ and the effective impurity operators $\hat{h}_{i}$ live in the extended MC effective Fock space,
~$\tilde{\mathcal{H}}^{\rm imp}_{\rm eff}=\mathcal{H}^{\rm imp}_{\rm eff}\otimes\mathcal{H}^{\rm bath}_{\rm 0}\;.$ (40)
As before, $\mathcal{H}^{\rm imp}_{\rm eff}$ spans the ground state manifold of $\hat{H}_{\rm bare}^{\rm imp}$ and $\mathcal{H}^{\rm bath}_{\rm 0}$ is the Fock space of the local bath orbital(s).
III.1 Corrections to the Kondo model
As the simplest example, we again consider the particle-hole symmetric AIM as our bare model, Eq. 24. The MC effective model can be found by following the above arguments. With a spin-$\tfrac{1}{2}$ operator $\hat{\mathbf{S}}_{d}$ representing the impurity and $\hat{c}_{0\sigma}$ for the local bath orbital, the effective impurity Fock space ~$\tilde{\mathcal{H}}^{\rm imp}_{\rm eff}$ is 8-dimensional. Since we have conserved spin $S$, spin projection $S_{z}$ and charge $Q$, we may label the basis states of the MC effective impurity Hamiltonian by these quantum numbers, $|Q,S;S_{z}\rangle$. Then $\tilde{H}^{\rm imp}_{\rm eff}(\boldsymbol{\theta})$ assumes a block diagonal structure, with each quantum number block in this case being simply a $1\times 1$ scalar. The allowed configurations are,
${}~{}Q~{}~{}$ ${}~{}S~{}~{}$ ${}~{}~{}S_{z}~{}~{}~{}$
±$\pm 1$ $\tfrac{1}{2}$$\pm\tfrac{1}{2}$
0 00 00
0 $1$ $0,~{}\pm 1$
where we specify the charge $Q$ with respect to the 2-electron half-filled state. Due to SU(2) spin symmetry, states with the same $Q$ and $S$ are degenerate; and due to particle-hole symmetry, states with $Q=\pm 1$ are degenerate. The most general MC effective Hamiltonian defined in this Fock space, consistent with these symmetries, can therefore be parameterized by just three distinct Hamiltonian matrix elements. Because in this case each quantum number block contains a unique basis state, the Hamiltonian operator can be entirely specified in terms of projectors onto these symmetry spaces,
$\hat{P}_{QS}=\sum_{S_{z}}|Q,S;S_{z}\rangle\langle Q,S;S_{z}|\;,$ (41)
according to,
$\tilde{H}^{\rm imp}_{\rm eff}(\boldsymbol{\theta})=\theta_{\pm 1,1/2}(\hat{P}_{-1,1/2}+\hat{P}_{+1,1/2})+\theta_{0,0}\hat{P}_{0,0}+\theta_{0,1}\hat{P}_{0,1}\;.$ (42)
These projectors can be written out explicitly as,